# Give and Take: An End-To-End Investigation of Giveaway Scam Conversion Rates

### v1

@@ 5.2. Twitter scam revenue
Table 2 summarizes the overall co-occurring revenue made across Twitter and YouTube, normalizing payment values across cryptocurrencies using the average USD price of each coin on the day in which the payment took place. [^7] We used historical data from https://finance.yahoo.com/crypto/. In total, we estimate that Twitter-based giveaway scams yielded \$2.7M in revenue. If we include all incoming transactions to these addresses then the total jumps to \$6.6M; as suggested above, the real revenue likely lies between the two numbers.
@@
Table 2. Revenue of giveaway scams on Twitter and YouTube livestreams, considering only co-occurring payments made in BTC, ETH, or XRP. For completeness, we also include all payments to the addresses, irrespective of what might explain the origin of the payment.
@@ 5.4. Victim behavior
The total revenue was dominated by a small number of transactions: the top 24 Twitter-related payments captured 50% of the value, while the top 164 payments captured 90% of the value. YouTube scams exhibited a similar phenomenon: 20 payments captured 50% of the value and 147 payments 90% of the value. These results suggest that the giveaway scam revenue model is significantly influenced by large transactions, where the goal for promoting the scam is to cast a wide net and, in doing so, capture a small fraction of victims who will give large amounts (analogous to the whale profit model for free-to-play mobile games (udonis-mobile-whales24, )). Such a model notably differs from the email spam ecosystem, for instance, in which the variance in individual customer payments is moderate and leads to a business model driven by order volume from new and repeat customers (pharmleaks:usesec12, ).
@@ 6. Discussion
6.1. Comparing scam revenue
Previous studies of pharmaceutical spam estimated a conversion rate of 0.00001%, with spam generating \$100 per interaction (kanich2008spamalytics, ). While our conversion rates are not directly comparable (e.g., tweets or livestreams vs. emails), they are markedly higher than email-based spam: 0.12% of tweets and 0.0039% of livestream views yielded a victim.
@@ 6.2. Interventions
A natural question for any scam is where bottlenecks exist that might be ripe for disruption. Here, one choice appears to be the payment origins of victims, as at least 58% of victims relied on centralized exchanges. Early detection of payments to scammers via these exchanges represents a sizable and durable protection (metamask, ): scammers likely cannot influence the exchange that victims use, as we suspect most victims have low comfort with using cryptocurrency directly. In addition, similar to what banks already implement, exchanges can consider implementing additional warnings when users attempt to make large transactions, which constitute a large portion of the revenue in our study. That said, detecting the cryptocurrency addresses engaging in fraud and informing these exchanges promptly remain challenging and require collaboration across multiple sectors.
@@ Appendix B Stream Retrieval Pilot Study
In this section, we provide more details on our pilot study conducted July 1-14, 2023. We start by discussing our attempt to identify scam livestreams on Twitch (Section B.1). Our search yielded no results, and thus we excluded Twitch from future experiments. Next, we detail how we chose the set keywords for searching on YouTube. Then, we go over the characteristics of a subset of the scam livestreams identified during the pilot study. We conclude with a short description of four anti-bot mechanisms employed by scam sites and the strategies we employed to counter them.
@@ B.1. Twitch details
The Twitch API allows callers to retrieve all livestreams at a given point in time. We retrieve the list of all livestreams every 30 minutes, which is roughly the amount of time we need to record and process the streams in that batch.
After retrieving all streams we identify the ones that are the most worth recording; i.e., that seem most likely to be promoting giveaway scams. The YouTube API already performs filtering, as it returns only streams related to the provided keywords. To filter out irrelevant streams on Twitch, we use their title, tags, and category. We start by marking a stream as relevant if its tags or title contain any relevant keywords. We derived our set of relevant keywords from the keywords used by CryptoScamTracker, but removed 16 non-crypto keywords (e.g., “event” and “give”) as they introduced too many irrelevant results. After filtering out streams that are not relevant, we then further remove any streams that are games using the category of the stream. This step effectively limits the number of streams to around 250.
@@
Twitch randomly inserts a 15-second advertisement clip before the actual stream content. To account for this issue, we record every Twitch stream for 20 seconds. For chat messages, Twitch’s API does not provide history for chat messages. As a result, we record the Twitch chat for the entire duration of the stream.
@@
Figure 5. Effectiveness of keywords
Category Keywords used
Coin names bitcoin btc ethereum eth tether usdt ripple xrp bnb “usd coin” usdc cardano ada coin ada dogecoin doge solana sol tron trx litecoin ltc polkadot dot polygon matic wrapped bitcoin wbtc bitcoin cash bch toncoin ton dai avalanche avax shiba inu shib binance usd busd algorand algo hex cryptocurrency crypto
Domain keywords kf event musk elon give coin shiba drop double get doge kefu vitalik claim binance hoskinson free charles star garling
HTML keywords giveaway participate send address rules crypto bonus immediately hurry
Table 3. Keywords used to find relevant streams.
B.2. Effectiveness of YouTube Search Keywords
Based on our pilot analysis, we opted for a set of keywords consisting of the names and ticker symbols of the top 20 cryptocurrencies (as listed on coinmarketcap.com on July 1, 2023), [^9] In three cases (for ADA, SOL, and DOT), we appended the word “coin” to the ticker symbol to avoid introducing irrelevant streams. along with two catch-all terms: “cryptocurrency” and “crypto.” We added other keywords used by the CryptoScamTracker tool (li2023giveaway, , Table XI) to increase coverage. We found that many of the additional keywords were not useful in expanding our list of relevant streams, and thus stopped expanding our keyword set.
In more detail, we examine the incremental contribution of each search keyword (Table 3) to the set of livestreams returned from the YouTube API. Specifically, we compute the number of YouTube streams (over the 14-day pilot study) that contain search keywords in their metadata (title and description). If one stream contains multiple keywords, we give credit evenly among all keywords.
Figure 5 shows the result. Overall, around 55% of the streams contained at least one of the search keywords, and the top 20 keywords account for 90% of those streams. Among the streams that do not contain any search keyword, roughly 50% are not in English. Sampling English streams that do not contain any keyword, we find that they are either on crypto-adjacent topics (e.g., trading) or completely irrelevant topics (e.g., live street camera).
B.3. Characteristics of scam livestreams
In our 14-day pilot study, we sought to understand the basic characteristics of scam livestreams and websites, which in turn informed the configuration of the parameters in our pipeline (e.g., for how long we record a scam livestream). Specifically, we are interested in two questions: (1) how long a QR code lasts once it appears and (2) how many chat messages are posted during the lifespan of a scam livestream. We started by running our pipeline at maximum capacity on YouTube, with the default setup of recording a stream for two seconds as well as retrieving up to 70 historical chat messages every 7.5 minutes. In addition, for a subset of the livestreams, we kept recording the stream upon detecting a QR code.
With this setup, we manually identified 276 streams that promoted 59 unique giveaway websites. To answer the first question, we examined how long a QR code lasted after we first detected it. For 41 scam livestreams, we continued to record them upon detecting a QR code. We found that for the majority of the livestreams, a QR code remained displayed continuously for a reasonable amount of time (mean 7,200 seconds and median 3,140 seconds), except for one case where the QR code was displayed periodically for around 15 seconds. This finding suggests that two-second samples are enough to capture a QR code in the livestream, if present. For the latter question, we examined scam streams that had chat messages. We found that all scam streams had few chat messages (less than 10) and no user interactions, and thus decided that 70 historical chat messages every 7.5 minutes is sufficient.
B.4. Cloaking Behaviors of Scam Websites
As part of our pilot, we identified four types of cloaking behavior: (1) IP-based cloaking, in which requests from an institutional network resulted in 403 responses but ones from a residential network did not; (2) user agent-based cloaking, in which requests from browsers that were not running on Windows or Mac resulted in 403 responses; (3) customized front pages, in which human interaction was required to click on a button or select a desired cryptocurrency; and (4) anti-bot detection through Cloudflare. As previously mentioned, to counter these behaviors, we (1) used a commercial VPN to mask our IP address; (2) spoofed our User-Agent string to appear as a popular device and browser; (3) added a heuristic module to click through common front pages; and (4) verified our crawler with Cloudflare to avoid triggering their anti-bot detection.

### v2

@@ 5.2. Twitter scam revenue
Table 2 summarizes the overall co-occurring revenue made across Twitter and YouTube, normalizing payment values across cryptocurrencies using the average USD price of each coin on the day of the payment. [^9] We used historical data from https://finance.yahoo.com/crypto/. In total, we estimate that Twitter-based giveaway scams yielded \$2.7M in revenue. If we include all incoming transactions to these addresses then the total jumps to \$6.6M; as suggested above, the real revenue likely lies between the two numbers.
@@
Table 2. Revenue of giveaway scams on Twitter and YouTube livestreams, considering only co-occurring payments made in BTC, ETH, or XRP. For completeness, we also include all payments to the addresses, irrespective of what might explain the payment origin.
@@ 5.4. Victim behavior
The total revenue was dominated by a small number of transactions: the top 24 Twitter-related payments captured 50% of the value, while the top 164 payments captured 90% of the value. YouTube scams exhibited a similar phenomenon: 20 payments captured 50% of the value and 147 payments 90% of the value. These results suggest that the giveaway scam revenue model is significantly influenced by large transactions, where the goal for promoting the scam is to cast a wide net and, in doing so, capture a small fraction of victims who will give large amounts (analogous to the whale profit model for free-to-play mobile games (udonis-mobile-whales24, )). Such a model notably differs from email spam, in which the variance in individual customer payments is moderate and leads to a business model driven by order volume from new and repeat customers (pharmleaks:usesec12, ).
@@ 6. Discussion
6.1. Approximating scam revenue
Previous studies of pharmaceutical spam estimated a conversion rate of 0.00001% (kanich2008spamalytics, ). While our conversion rates are not directly comparable (e.g., tweets or livestreams vs. emails), they are markedly higher than email-based spam: 0.12% of tweets and 0.0039% of livestream views yielded a victim. We again acknowledge that the conversion rate for Twitter is based on the number of tweets rather than the number of views (as the number of views per tweet is unknown), the set of payments might include irrelevant or exclude relevant transactions (thus artificially inflating or deflating the derived rate), and the analyses take place in different time periods. Nevertheless, these factors are unlikely to account for the orders-of-magnitude difference with email spam.
@@ 6.2. Interventions
A natural question for any scam is where bottlenecks exist that might be ripe for disruption. Here, one choice appears to be the payment origins of victims, as at least 58% of victims relied on centralized exchanges. Early detection of payments to scammers via these exchanges represents a sizable and durable protection (metamask, ): scammers likely cannot influence the exchange that victims use, as we suspect most victims have low comfort with using cryptocurrency directly. In addition, similar to what banks already implement, exchanges could implement additional warnings when users attempt to make large transactions, which constitute a large portion of the revenue in our study. That said, detecting the cryptocurrency addresses engaging in fraud and informing these exchanges promptly remain challenging problems and require collaboration across multiple sectors.
@@ 7. Conclusions
Acknowledgements
We thank our anonymous shepherd and reviewers for their insightful suggestions and feedback. We also thank Xigao Li for his advice on setting up the crawling infrastructure, Babak (Bob) Amin Azad for helping us verify our bot with Cloudflare, and Jennifer Folkestad and Cindy Moore for operational support. This work was supported in part by National Science Foundation grant CNS-2152644.
@@ Appendix B Stream Retrieval Pilot Study
In this section, we provide more details on our pilot study conducted July 1-14, 2023. We start by discussing our attempt to identify scam livestreams on Twitch (Section B.1). Our search yielded no results, and thus we excluded Twitch from future experiments. We then describe how the various keywords we used for searching for livestreams on YouTube incrementally contributed to the search results.
@@ B.1. Twitch details
The Twitch API allows callers to retrieve all livestreams at a given point in time. We retrieved the list of all livestreams every 30 minutes, which is roughly the amount of time we needed to record and process the streams in that batch.
After retrieving all streams we identified the ones that are the most worth recording; i.e., that seem most likely to be promoting giveaway scams. (The YouTube API already performs filtering, as it returns only streams related to the provided keywords.) To filter out irrelevant streams on Twitch, we used their title, tags, and category. We started by marking a stream as relevant if its tags or title contain any relevant keywords. We derived our set of relevant keywords from the keywords used by CryptoScamTracker, but removed 16 non-crypto keywords (e.g., “event” and “give”) as they introduced too many irrelevant results. After filtering out streams that are not relevant, we then further removed any streams that are games using the category of the stream. This step effectively limits the number of streams to around 250.
@@
Twitch randomly inserts a 15-second advertisement clip before the actual stream content. To account for this issue, we recorded every Twitch stream for 20 seconds. For chat messages, Twitch’s API does not provide history for chat messages. As a result, we recorded the Twitch chat for the entire duration of the stream.
@@
B.2. Keyword details
We examine the incremental contribution of each search keyword (Table 3) to the set of livestreams returned from the YouTube API by computing the number of YouTube streams (over the 14-day pilot study) that contain search keywords in their metadata (title and description). If one stream contains multiple keywords, we give credit evenly among all keywords. Figure 5 shows the result. Overall, around 55% of the streams contained at least one of the search keywords, and the top 20 keywords account for 90% of those streams. Among the streams that do not contain any search keyword, roughly 50% are not in English. Sampling English streams that do not contain any keyword, we find that they are either on crypto-adjacent topics (e.g., trading) or completely irrelevant topics (e.g., a live street camera).